OpenStudy (anonymous):

Will fan and medal if you answer right The rate law for a hypothetical reaction is rate = k [A]2. When the concentration is 0.10 moles/liter, the rate is 2.7 × 10-5 M*s-1. What is the value of k? 2.7 × 10-3 2.7 × 10-3 M-1s-1 7.2 × 105 M-1s-1 2.7 × 10-4 M-1s-1

Vocaloid (vocaloid):

plug in 2.7 * 10^-5 for rate, and 0.1 for A, then solve for k

OpenStudy (anonymous):

I got 2.7 8 10^-6

OpenStudy (anonymous):

Which is not even in the answer choice

Vocaloid (vocaloid):

check your math again...

Vocaloid (vocaloid):

(2.7*10^-5)/(0.1^2) = ?

OpenStudy (anonymous):

No Clue. I'm NOT this advanced in math yet

OpenStudy (anonymous):

oki let me try

OpenStudy (anonymous):

i got .0027, idk anymore

Vocaloid (vocaloid):

right, 0.0027, now convert to scientific notation

OpenStudy (anonymous):

how do you convert to scientific notation on calculator

Vocaloid (vocaloid):

uh, I'm not sure you can do that on a calculator 0.0027 = 2.7 * 10 ^ x find x

OpenStudy (anonymous):

10^-4

OpenStudy (anonymous):

?

Vocaloid (vocaloid):

not quite, try again... how many places do you have to move the decimal point to get from 0.0027 to 2.7?

OpenStudy (anonymous):

3

Vocaloid (vocaloid):

right! so, 0.0027 = 2.7 * 10^(-3) now, the units of k are M-1s-1 so our answer becomes 2.7 * 10^(-3) M-1s-1

OpenStudy (anonymous):

Thank you